# Cryptocurrency as an Investable Asset Class: Coming of Age

### v2

@@ 5 CONCLUSIONS
In the 16 years since Nakamoto, cryptocurrencies have reached a certain stage of maturity: they can now be studied using the same measurement, data quality, and research rigor applied to other asset classes. The market is now worth more than four trillion dollars. Major financial firms are using cryptocurrency in their businesses.
We show that there are important similarities between cryptocurrency markets and traditional asset markets—Sharpe ratios are broadly comparable, and a small number of factors and their higher orders summarize the cross-section of returns. On the other hand, cryptocurrency also has its own character—frequent jumps and blockchain information driving prices. The cryptocurrency market has become more mature and is emerging as an asset class with its own distinct characteristics that still, however, needs more data, research, supervision, enforcement, and a better legal and regulatory framework.
Given the maturation of cryptocurrency into an asset class, this review shows that cryptocurrency should be studied the same way we study any asset. The right way to understand it is exactly through traditional finance tools—empirical asset pricing. The ten facts we present summarize the key takeaways from this view.

### v3

@@ 5 CONCLUSIONS
In the 17 years since Nakamoto, cryptocurrencies have reached a certain stage of maturity: they can now be studied using the same measurement, data quality, and research rigor applied to other asset classes. By the end of our sample period (September 2025), aggregate cryptocurrency market capitalization is about \$4 trillion. Major financial firms are using cryptocurrency in their businesses.
We show that there are important similarities between cryptocurrency markets and traditional asset markets—Sharpe ratios so far are broadly comparable, and a small number of factors and their higher orders summarize the cross-section of returns. On the other hand, cryptocurrency also has its own character—frequent jumps and blockchain information driving prices. The cryptocurrency market has become more mature and is emerging as an asset class with its own distinct characteristics that still, however, needs more data, research, supervision, enforcement, and a better legal and regulatory framework.
Given the maturation of cryptocurrency into an asset class, this review shows that cryptocurrency should be studied the same way we study any asset. The right way to understand it is exactly through traditional finance tools—empirical asset pricing. The seven stylized facts we present summarize the key takeaways from this view.